Consider the following observed values: (-5,-2) (-3,1) (0,4) (2,6) (1,3)

(1) find the estimated regression line based on the observed data

(2) for each xi, compute fitted value of yi

(3) compute the residuals ei

(4) calculate the coefficient of determination
Answer #1

1)The regression line is given as y=3.4+x

It is obtained by the normal equations from the method of least squares

2)Predicted value of y is obtained by substituting the values of x in the equation given in 1.

3)Residual ,ei= actual y - predicted y

Direct Marketing Graphs Utlities Add-ons : PRE 1 PRE 1 X RES 1 var 1 -5.00 -2.00 -1.60000 -.40000 2 -3.00 1.00 40000 60000 004) Coefficient of determination,R2= MSS/TSS = 95.97/105 = 0.914